OpenStudy (anonymous):

What is the equation of a circle with a center at (-4, -8) and a radius of 1.9?

OpenStudy (anonymous):

use equation of circle again http://www.regentsprep.org/Regents/math/algtrig/ATC1/circlelesson.htm

OpenStudy (anonymous):

(x-k)^2+(y-h)^2=(r)^2

OpenStudy (anonymous):

just plug in (x+4)^2+(y+8)^2=(1.9)^2

OpenStudy (anonymous):

(x+16)+(y+64)=3.61 ?

OpenStudy (anonymous):

no silly you only square the r side :)

OpenStudy (anonymous):

(x+4)^2+(y+8)^2=(1.9)^2 (x+4)^2+(y+8)^2=3.61

OpenStudy (anonymous):

oopsy lol thanks!

OpenStudy (anonymous):

:) btw for example (x+4)^2 is not x+16 the square is not distributed to the 4. (x+4)^2 means (x+4)*(x+4). usually they dont require you to expand it. But to do it you must use the FOIL. However you dont have to.